OpenStudy (shaik0124):

a man travels from a to b at 5 kmph and reaches b 20 minutes late .if his speed had been 7.5kmph he would have reached his office 12 minutes early find the distance between a to b

OpenStudy (shaik0124):

b=office

OpenStudy (jack1):

increase in 2.5 mph = reduction of 32 minutes in driving time... does this help?

OpenStudy (shaik0124):

how?

OpenStudy (jack1):

hint - convert all to one unit (so miles per minute, or convert 32 minutes into a decimal of how many hours that is)

OpenStudy (rational):

do we have two equations d = 5*(t+20) d = 7.5*(t-12)

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} d=5(t+\dfrac{20}{60})=7.5(t-\dfrac{12}{60})\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (jack1):

key formula: \(\huge \frac {miles~ travelled}{speed} = time~of ~travel\) journey 1 \(\large x~ miles \times \frac {1~ hour}{5~ miles} = y~ hours\) journey 2 \(\large x ~miles \times \frac {1~ hour}{7.5 ~miles} = (y-0.2)~ hours\) rearrange equation 1 in terms of x, and substitute into equation 2 journey 1 \(\large x~ miles \times \frac {1~ hour}{5~ miles} = y~ hours\) \(\large x = 5y \) sub x = 5y into journey 2 equation journey 2 \(\large x~ miles \times \frac {1 ~hour}{7.5 ~miles} = (y-0.2) hours\) \(\large 5y~ \times \frac {1 }{7.5} = (y-0.2)~ hours\) you now have only one variable in this equation, can you solve this and find y (which was the time taken for the original trip). Then use this to solve the distance.

OpenStudy (jack1):

bugger... the 0.2 is incorrect, that's 12 minutes in hours... it should be 32 minutes difference, sooooo 8/15 hours, not 2/10 hours... sorry

OpenStudy (jack1):

\(\large 5y~ \times \frac {1 }{7.5} = (y-(\frac 8{15})~ hours\)